OpenStudy (amber15):

User: which numbers are divisible by 6? 62 ,248,930, 124,310,558,155 ,7812,93,31,465,279

OpenStudy (mertsj):

The ones that are divisible by 2 and by 3

OpenStudy (mertsj):

In other words if the sum of the digits is a multiple of 3 and if it is even.

OpenStudy (amber15):

ok

OpenStudy (anonymous):

930, 558,7812

OpenStudy (amber15):

thank you

OpenStudy (anonymous):

It's very fast using Mathematica. Copy and Paste, type /6 and then press enter on the numeric pad. {62,248,930,124,310,558,155,7812,93,31,465,279}/6\[\left\{\frac{31}{3},\frac{124}{3},155,\frac{62}{3},\frac{155}{3},93,\frac{155}{6},1302,\frac{31}{2},\frac{31}{6},\frac{155}{2},\frac{93}{2}\right\} \]

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (amber15):

your welcome

OpenStudy (anonymous):

The answer is 930, 558 and 7812